completed application cost each= $ 81

out of all of the applications completed, 80% are approved and 67% funded.

Each application completed converts to 1*0.8*0.67 = 0.536 customers

Cost per customer = $81/0.536 = $151

Promotional costs = $100

Effective Acquisition cost = $251

Retention rate for online customers=55%

CLV formula = mr/(1+i-r) – Acquisition Cost

CLV for online customers = 300*(0.55)/(1-0.55) -$251 = $117

Given the information above and keeping everything constant, at what retention rate will the CLV(customer lifetime value) for an online customer be 0.
Answer #1

CLV = 300*r/(1+i-r) - 251

As per given equation, of CLV for online customers, i = 0

Equating the above CLV equation to 0, we get,

CLV = 300*r/(1-r) - 251 = 0

or, r/(1-r) = 251/300

or, r = 0.8367*(1-r)

or, r*1.8367 = 0.8367

or, r = 0.8367/1.8367

r = 0.4555

CLV for an online customer will be 0, for retention rate equal to 45.55 %